OpenStudy (anonymous):

i'm suppose to find the equation of a line i only know the x and y intercept x being -1 and y being -5. i need them in slop intercept form but confused as to how to find it

OpenStudy (saifoo.khan):

u just have to find the slope thats all.

OpenStudy (saifoo.khan):

u have two cordinates right? (-1,0) & (0, -5) use these to find the slope.

OpenStudy (saifoo.khan):

Slope = -5/1 = -5

OpenStudy (saifoo.khan):

Now, put these in the equation. y = mx + b y = -5x -5

OpenStudy (anonymous):

all i have on my paper is x-incept= -1, y intercept=-5 i dont have any coordinates

OpenStudy (saifoo.khan):

when x intercpts, y = 0 when y intercpts x = 0 so, make the cordinates by youself.

OpenStudy (saifoo.khan):

x = (-1,0) & y= (0, -5)

OpenStudy (anonymous):

so then i find the slope with those right

OpenStudy (saifoo.khan):

Right.

OpenStudy (anonymous):

so it would be 5 over 1?

OpenStudy (saifoo.khan):

\[m = \frac{-5-0}{0-(-1)}\]

OpenStudy (saifoo.khan):

\[m = \frac{-5}{1}\]

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

ok i got is so then how would i put it in slop-intercept form?

OpenStudy (saifoo.khan):

u got the y intercpt right?

OpenStudy (saifoo.khan):

and u got the slope too.

OpenStudy (anonymous):

yea

OpenStudy (saifoo.khan):

now u also know that in an equation: y = mx + b where m is slope and b is the y-intercpt.. just simply insert.

OpenStudy (anonymous):

so it should look like y=-5(-1)-5?

OpenStudy (saifoo.khan):

why u wrote -1?

OpenStudy (saifoo.khan):

it isnt needed. :p

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so y=-5x-5?

OpenStudy (saifoo.khan):

right.

OpenStudy (anonymous):

yay thank you so much(:

OpenStudy (saifoo.khan):

;] you welcome soo much! :]

OpenStudy (anonymous):

i feel really dum cuz the last time i did this was 2 month ago and i dont remember anything :/

OpenStudy (saifoo.khan):

LOL, np.. i think now u have recovered some of it. ^.^

OpenStudy (anonymous):

i hope cuz i have a lot more problems ^^

OpenStudy (saifoo.khan):

np, if u have any problem.. v are here! :)

OpenStudy (anonymous):

thanks im really glad i found this place(:

OpenStudy (saifoo.khan):

JK.

OpenStudy (anonymous):

mostly? lol

OpenStudy (saifoo.khan):

sorry, i mean here totally.

OpenStudy (saifoo.khan):

LOL.

hero (hero):

Some people get full explanations, others just get answers....Some people get free medals...Some people get lots of free medals....Some just get ALL the medals.....for FREE!

OpenStudy (anonymous):

lol oh ok

OpenStudy (saifoo.khan):

HAHA, those some ppl do give explaination too!

OpenStudy (saifoo.khan):

while other some ppl get Jealous, JK JK

OpenStudy (anonymous):

lol wow

OpenStudy (saifoo.khan):

\[\Huge HAHA!\]

hero (hero):

Some people earned they're guru, other people were given guru

hero (hero):

\Huge Huge Fake

hero (hero):

damn it!

OpenStudy (saifoo.khan):

how can some ppl be GIVEN GURU?

OpenStudy (saifoo.khan):

\[HAHA!\]

hero (hero):

Because people give it to them

hero (hero):

And they give it to the one who does nothing but give answers all dayu

hero (hero):

Everything on here is free...including crime. Crime is free too

hero (hero):

Oh wait...

hero (hero):

That's like an oxymoron on something, there's no such thing as free crime

OpenStudy (saifoo.khan):

LOL, WHT?

OpenStudy (anonymous):

im lost o.O

hero (hero):

Saif...you're practicing the art of free crime....it's a mafia tradition

hero (hero):

Don't worry, I won't tell anybody

OpenStudy (saifoo.khan):

HAHA, Shhhhhhhhhhh. ;)